# Superficial Self-Improved Reasoners Benefit from Model Merging

### v1

@@ 5.1 Setup
We train the model on MATH Hendrycks et al. (2021) and GSM-8K Cobbe et al. (2021) datasets correspondingly to evaluate the in domain reasoning ability of the model, while evaluate it on MAWPS Koncel-Kedziorski et al. (2016), SAT-Math Zhong et al. (2024) datasets to evaluate the out-of-domain reasoning ability.
@@
We include three LLMs at different scales (Qwen2.5-0.5B-Instruct, Qwen2.5-1.5B-Instruct Yang and et al. (2024) and Llama2-7B Touvron et al. (2023)) for self-improvement training. For the distillation experiment we include stronger teacher models Qwen2.5-7B-Instruct for distillation.
@@
We evaluate our method by comparing it with four baselines. First, we consider Vanilla (STaR Zelikman et al. (2022)) , which iteratively generates reasoning data following the procedure in Section 3 for self-improvement. Second, we include Data Mixture Shumailov et al. (2023), which mitigates performance degradation by mixing a portion of data from previous iterations. Third, we compare with Data Accumulation Gerstgrasser et al. (2024), which demonstrates that accumulating synthetic data across iterations can prevent model collapse.
@@
[…]
We also provide the training and implementation details in Appendix B.
5.2 Main Results for Self-improvement
To answer the first research question, we conducted extensive experiments in a model collapse setting (iterative self-improvement) using two mathematical reasoning datasets, GSM8K and MATH. The results, shown in Figure 6, highlight that across three self-improvement iterations with three different LLMs, model collapse occurs in the first or second iteration for the baseline methods. In contrast, our method successfully avoids model collapse and achieves the best performance after applying model merging. Not only does our method significantly delay model collapse, but it also maintains superior performance across all iterations. Moreover, we observe that LLMs of all scales benefit from our model merging strategy, with smaller models suffering more severely from model collapse in the absence of this approach. Given the rising importance of test-time computing Snell et al. (2024), we further evaluate our method by generating multiple answers and measuring pass@$k$ accuracy. As shown in Figure 8, our method consistently improves performance as k increases and outperforms both the base models and the SFT models.
@@ 5.3 OOD Generalization Results
To answer the second research question, We evaluate the checkpoints from Section 5.2 using two out-of-domain (OOD) math reasoning datasets: SAT Math and MAWPS. The results, presented in Figure 7, show that while all other baselines suffer significant OOD performance degradation after iterative self-improvement, our method consistently restores performance after each model merging step and, in some cases, even surpasses the original base model. The only exception is the Qwen2.5-0.5B-Instruct model on the MAWPS dataset. We hypothesize that this dataset closely resembles the in-domain data, where extensive ID training significantly improves performance (from 12.8% to 62.9%), which may cause a slight degradation during the model merging process. Overall, these results demonstrate the great potential of our method, as it successfully mitigates the generalization drop commonly observed during SFT on specific tasks or datasets. Additional OOD datasets results can be found in Appendix F.
@@ 5.4 Distillation from Stronger Models
Considering self-improvement may only one of paradigm for LLM distillation, we extend our method to a broader field to answer the third research question. We distill stronger Qwen2.5-7B-Instruction model into the weaker Qwen2.5-1.5B-Instruction and Llama-2-7B models. The results in Table 2 demonstrate that the merged models consistently improve or maintain comparable performance on ID tasks, while often achieving significant improvements in OOD performance. This indicates that our method not only preserves task-specific performance but also enhances the model’s generalized reasoning ability when distilling from a stronger teacher model.
@@ 6 Conclusion
This study identifies that self-improved LLM reasoners still lack generalized reasoning capability on OOD datasets, although they have better ID reasoning performances. We refer to this phenomenon as Superficial Self-Improved Reasoners and investigate its underlying causes. Our analysis reveals that during the self-improvement process, the layer weights change order doesn’t match the order of importance layers. This mismatch suggests that instead of solely learning to reason, the model also memorizes the training data. To address this issue, we propose Iterative Model Merging and extensive experiments demonstrate the effectiveness of our method: it not only mitigates model collapse but also have a generalized reasoning capability.
@@
The proposed Iterative Model Merging (IMM) method currently employs a fixed-weight merging mechanism between the original and self-improved models. However, more advanced strategies, such as dynamic or layer-adaptive merging, could provide further improvements. Additionally, although IMM has proven to effectively maintain generalized reasoning capabilities, it doesn’t investigate the strategy of mixing real and synthetic data appropriately, which could further enhance the trade-offs between reasoning improvement and generalization. We leave the exploration of advanced merging mechanisms and the optimal mixture ratio of real and synthetic data for future work.
Appendix A Chain of Thought Prompting for Data Synthesize
We use chain-of-thought prompting Wei et al. (2022) to generate answers. For MATH and GSM8K datasets, we both give 10 examples in the instructions for in-context generation. The prompting examples are given in Table 4.
Appendix B Training and implementation Details
[…]
@@ Appendix B Training and implementation Details
Appendix C Superficial Reasoning Finetuning Exists When Real Data is Limited
We also find that even using real but limited data, Superficial Reasoning Synthetic Finetuning still exists. As Figure 9 shows, the middle layers change most compared with the early and late layers, while Figure 3 already shows early and late layers are more important for reasoning. However, utilizing real data prevents the model from overfitting itself by using self-generated data. This is also verified by Figure 9: model’s reasoning layer (early and late layers) changed more (learn more reasoning capability) when training with real data, reasoning-trivial layers (middle layers)’s weight change is close to middle layers when training with synthetic data.
@@ Appendix C Superficial Reasoning Finetuning Exists When Real Data is Limited
Appendix D A bitter Lesson: Not All LLMs Can Self-improve
During our experiment, we also find that not all the LLMs can self-improve on reasoning tasks. If LLM’s performance decreases after SFT, then our method may not let the merged model have a better performance compared with the original model and the model after SFT. This usually happens when the original model already has a good performance (reasoning ability), and learned reasoning ability can’t offset the generalization loss. Unlike previous works that improve generalization and robustness via random smoothing Cohen et al. (2019); Yuan et al. (2024c) or regularization Virmaux and Scaman (2018); Ye et al. (2023); Yuan et al. (2024b), we utilize model merge to improve generalization and mitigate superficial self-improved reasoners phenomenon.
Appendix E The Connection to Catastrophic Forgetting
@@
Appendix F Additional Experiments
Here we provide the additional experiment results for evaluating the layer importance for Qwen2.5-1.5B base model on reasoning datasets MATH. Similar to stronger reasoning model Qwen2.5-1.5B-Math, the importance layer for reasoning is early and late layers, as demonstrated in Figure 10. We also provide OOD performance on additional datasets SVAMP Patel et al. (2021), ASDiv Miao et al. (2020), MathQA Amini et al. (2019) and MMLU-stem Hendrycks et al. (2020). IMM keeps the OOD reasoning capability as shown in Table 3.

### v2

@@ 5.1 Setup
We train the model on MATH Hendrycks et al. (2021) and GSM-8K Cobbe et al. (2021) datasets correspondingly to evaluate the in-domain reasoning ability of the model, while evaluate it on MAWPS Koncel-Kedziorski et al. (2016), SAT-Math Zhong et al. (2024) datasets to evaluate the out-of-domain reasoning ability.
@@
We include three LLMs at different scales (Qwen2.5-0.5B-Instruct, Qwen2.5-1.5B-Instruct Yang and et al. (2024) and Llama2-7B Touvron et al. (2023)) for self-improvement training. For the distillation experiments, we include stronger teacher models Qwen2.5-7B-Instruct for distillation. We also provide the recent model Llama3-8B performance in Appendix B.8
@@
We evaluate our method by comparing it with four baselines. First, we consider Vanilla (STaR Zelikman et al. (2022)) , which iteratively generates reasoning data following the procedure in Section 3 for self-improvement. Second, we include Data Mixture Shumailov et al. (2023), which mitigates performance degradation by mixing a portion of data from previous iterations. Third, we compare with Data Accumulation Gerstgrasser et al. (2024), which demonstrates that accumulating synthetic data across iterations can prevent model collapse. We also provide a comparison of SFT interventions in Appendix B.4.
@@
We evaluate the model performance by computing $\text{pass@}k=\mathbb{E}_{\mathcal{D}_{G}}\left[1-\frac{\binom{M-c}{k}}{\binom{M}{k}}\right$], where $c$ is the number of correct answers, out of total answer $M$ and $\mathbb{E}_{\mathcal{D}_{G}}[\cdot]$ is the expectation for overall generated dataset $\mathcal{D}_{G}$. Therefore, $\text{pass@}k$ measures the fraction of unique questions that have at least one correct answer when sampling $k$ answers per question from the model.
Additional training and implementation details are provided in Appendix A.2.
5.2 ID Results with Self-improvement
To answer research question (i), we conducted extensive experiments in a model collapse setting (iterative self-improvement) using two mathematical reasoning datasets, GSM8K and MATH. The results, shown in Figure 6, highlight that across three self-improvement iterations with three different LLMs, model collapse occurs in the first or second iteration for the baseline methods. In contrast, our method successfully avoids model collapse and achieves the best performance after applying model merging. Not only does our method significantly delay model collapse, but it also maintains superior performance across all iterations. Moreover, we observe that LLMs of all scales benefit from our model merging strategy, with smaller models suffering more severely from model collapse in the absence of this approach. Given the rising importance of test-time computing Snell et al. (2025), we further evaluate our method by generating multiple answers and measuring pass@$k$ accuracy. As shown in Figure 8 (more results are presented in Appendix B.7), our method consistently improves performance as k increases and outperforms both the base models and the SFT models.
@@ 5.3 OOD Generalization Results
To answer research question (ii), We evaluate the checkpoints from Section 5.2 using OOD math reasoning datasets: SAT Math and MAWPS. Additional OOD datasets results can be found in Appendix B.3. The results, presented in Figure 7, show that while all other baselines suffer significant OOD performance degradation after iterative self-improvement, our method consistently restores performance after each model merging step and, in some cases, even surpasses the original base model. The only exception is the Qwen2.5-0.5B-Instruct model on the MAWPS dataset. We hypothesize that this dataset closely resembles the in-domain data, where extensive ID training significantly improves performance, which causes a degradation during IMM. We further analyze this unexpected behavior in Appendix B.6. Overall, these results demonstrate the great potential of our method, as it successfully mitigates the generalization drop commonly observed during SFT.
@@ 5.4 Distillation from Stronger Models
Considering self-improvement may be only one of paradigms for LLM distillation, we extend our method to a broader field to answer research question (iii). We distill a stronger Qwen2.5-7B-Instruct model into the weaker Qwen2.5-1.5B-Instruct and Llama-2-7B models. The results in Table 2 demonstrate that IMM consistently improves or maintains comparable performance on ID tasks, while often achieving significant improvements in OOD performance. This indicates that IMM only preserves task-specific performance but also enhances the model’s generalized reasoning ability when distilling from the teacher model.
@@ 6 Conclusion
This study identifies that self-improved LLM reasoners still have the model collapse risk and lack generalized reasoning capability on OOD datasets. Our analysis reveals that the weight changes of layers doesn’t match the layer importance. This mismatch suggests that instead of solely learning to reason, the model also memorizes the training data. To address this issue, we propose the Iterative Model Merge and extensive experiments demonstrate the effectiveness of our method: it not only mitigates model collapse but also make model have generalized reasoning capability.
@@
The proposed Iterative Model Merging (IMM) method currently employs a fixed-weight merging mechanism between the original and self-improved models. However, more advanced strategies, such as dynamic or layer-adaptive merging, could provide further improvements. Additionally, although IMM has proven to be effective in maintaining generalized reasoning capabilities, it doesn’t investigate the strategy of mixing real and synthetic data appropriately, which could further enhance the trade-offs between reasoning improvement and generalization. We leave the exploration of advanced merging mechanisms and the optimal mixture ratio of real and synthetic data for future work.
7 Acknowledgement
This material is based upon work supported by the National Science Foundation under grant No. 2229876 and is supported in part by funds provided by the National Science Foundation, by the Department of Homeland Security, and by IBM. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or its federal agency and industry partners.
Appendix A Training and Implementation Details
A.1 Chain of Thought Prompting for Data Synthesis
We use chain-of-thought prompting Wei et al. (2022) to generate answers. For MATH and GSM8K datasets, we both give 10 examples in the instructions for in-context generation. The prompting examples are given in Table 11. We generate 3 candidate answers for GSM8K and 6 candidate answers for MATH to have comparable numbers of right answers.
A.2 Training Details
We use NVIDIA RTX 8 $\times$ A6000 to train the model with DeepSpeed Rajbhandari et al. (2020) distributed training framework. The number of training epoch is 3 and per device training batch size is 4. The gradient accumulation steps are set to 4 and the learning rate is 2e-5. The warm-up rate is 0.03. We use mixed precision training with bf16. We use DeepSpeed to distribute supervised fine-tuning model with ZeRO3, which partitions all three model states. We also use the vLLM library Kwon et al. (2023) to generate synthetic reasoning data with sampling temperature {0.2, 0.4, 0.6} to balance the diversity and accuracy of generated answers. Note that we use all models, data, and training tools solely for research purpose, which are consistent with their intended use.
The Model merging parameter in Section 3 is set to 0.5 to balance the base model and self-improved model. We use the setting in Section 5.4 to do the parameter analysis for $\alpha$. Table 3 shows that $\alpha=0.5$ can achieve a good balance between ID and OOD performance.
$\alpha$ 0.1 0.3 0.4 0.5 0.6 0.7 0.9
GSM8K 3.7 24.5 32.4 38.8 40.4 43.3 48.3
MATH 3.7 8.4 10.8 12.5 12.5 11.8 10.3
SAT_Math 25.3 27.4 27.7 28.1 26.7 24.8 18.8
MAWPS 64.4 69.0 76.8 76.6 69.3 64.3 57.2
Table 3: The parameter analysis for $\alpha$.
@@ A.2 Training Details
Datasets GSM8K MATH SAT_Math MAWPS
Vanilla SFT 58.5 32.5 50.0 85.9
Gradient-decay ($\gamma$=0.9) 59.2 32.8 53.8 84.2
Gradient-clipping (max_norm=2.0) 58.7 31.7 52.3 84.7
Weight-masking (TopP=0.3) 60.2 34.5 56.2 87.0
IMM 69.3 34.0 68.8 89.4
Table 5: Qwen2.5-1.5B-Instruct performance compared with SFT interventions in the first iteration.
Appendix B Additional Experiments and Analysis
B.1 Superficial Reasoning Finetuning Exists When Real Data Is Limited
We also find that even using real but limited data, Superficial Reasoning Synthetic Finetuning still exists. As Figure 9 shows, the middle layers change most compared with the early and late layers, while Figure 3 already shows that early and late layers are more important for reasoning. However, utilizing real data prevents the model from overfitting itself by using self-generated data. This is also verified by Figure 9: the model’s reasoning layer (early and late layers) changed more (learn more reasoning capability) when training with real data, the reasoning-trivial layers (middle layers)’s weight change is close to middle layers when training with synthetic data.
@@ B.1 Superficial Reasoning Finetuning Exists When Real Data Is Limited
B.2 Layer Importance
@@
Figure 11: The layer importance score for Qwen2.5-1.5B base model on reasoning dataset MBPP.
Here we provide the additional experiment results for evaluating the layer importance for Qwen2.5-1.5B base model on reasoning datasets MATH. Similar to stronger reasoning model Qwen2.5-1.5B-Math, the importance layer for reasoning is early and late layers, as demonstrated in Figure 10. We also observed similar behavior in other complex reasoning task code generation MBPP, as demonstrated in Figure 11.
B.3 OOD Performance
We also provide OOD performance on additional datasets SVAMP Patel et al. (2021), ASDiv Miao et al. (2020), MathQA Amini et al. (2019) and MMLU-stem Hendrycks et al. (2020). IMM keeps the OOD reasoning capability as shown in Table 4.
B.4 Comparison with SFT Interventions
We provide experimental results to compare these alternative interventions with IMM. As shown in Table 5, these methods generally do not outperform IMM, and in some cases are even outperformed by vanilla SFT. Compared with these interventions during SFT, IMM not only mitigate the overfitting reasoning finetuning, but also improve generalized reasoning capability through ensemble model merging. Our method is orthogonal to interventions for SFT, and provides a simple yet effective method to solve superficial self-improved reasoners phenomenon identified by this research.
B.5 Importance-based Weight Merge
Datasets GSM8K MATH SAT_Math MAWPS
I-IMM 44.2 27.5 49.3 2.1
IMM 44.2 27.4 56.2 3.4
Table 6: Comparison of IIMM and IMM across ID and OOD datasets.
We also experimented with weighting the merge ratio $\alpha$ per layer using the importance score $I$ defined in Eq. (3). As shown in Table 6, this approach occasionally improves in-domain (ID) performance but often performs worse on out-of-domain (OOD) datasets. We hypothesize that this is because weighting the merging process based on ID-specific importance scores leads to overfitting to the ID data, thereby sacrificing the model’s generalized reasoning capabilities on OOD tasks. Additionally, imbalanced merging rates across layers may introduce instability: when different layers are merged to varying degrees, the model can become internally inconsistent. In an extreme case, if some layers remain largely as base model layers while others are heavily adapted via SFT, this imbalance can degrade performance, as the layers are no longer "on the same page". We consider our merging method as a new way to increase the generalization of supervised learning, like other methods such as regularization Jin et al. (2024) and meta-tuning Guo et al. (2024a).
B.6 Analysis on Unexpected Behavior
Model Base SFT Merge
[…]
Dataset Base SFT Merge
[…]
MathQA $5.5 \ul62$.0
[…]
OOD performance drops for Qwen2.5-1.5B on MAWPS dataset, and here we conduct more experiments to analyze this behavior. We found that (Table 7) small models (e.g., 0.5B and 1.5B) only suffer significant performance degradation on the MAWPS dataset after model merging. In contrast, larger models (e.g., 7B) achieve the best performance on MAWPS, benefiting more from IMM. Despite this drop on MAWPS, smaller models still show performance improvements on other OOD datasets. For instance, Table 8 shows that the 1.5B model outperforms both the Base and SFT versions on 5 OOD datasets. Therefore, we attribute the performance degradation on MAWPS primarily to two factors: (1) potential distributional differences in MAWPS compared to other datasets, and (2) the limited parameter capacity of small models, which may lack sufficient redundancy to support robust merging without trade-offs.
B.7 Additional Test-time Computing Results
We evaluate our method by generating multiple answers and measuring pass@$k$ accuracy for MATH dataset. As shown in Figure 13, our method consistently improves performance as k increases and outperforms both the base models and the SFT models.
B.8 IMM with the Recent Model
Table 9 shows that for Llama3-8B model, IMM improves the ID performance and keeps comparable OOD performance, while vanilla SFT suffers from model collapse in ID datasets and severe degradation on OOD datasets.
Figure 12: The percentage of the weight change over layers for finetuning Qwen2.5-1.5B in different iterations.
B.9 Weight Change for Different Iterations
After the first model weight merge, the parameter updates for the layers critical for reasoning still remain minimal, as illustrated in Figure 4. However, we continue to analyze the weight change across different layers and find that, although IMM uses an average merge rate across different layers, it improves the model’s generalized reasoning capability, which makes the weight of reasoning-critical layers change more in the next iterations. Figure 12 shows that, in the next iterations, the reasoning-critical layers (early and late layers) change more weight change compared with the reasoning-trivial layers (middle layers), indicating the model learns the generalized reasoning capability after IMM. Also, although IMM uses a uniform merge rate $\alpha$ across all layers, the absolute weight change difference between reasoning-critical layers and reasoning-trivial layers becomes smaller compared with SFT. This small difference accumulates over the course of the iterative self-improvement process. As a result, IMM achieves a relatively more balanced distribution of weight changes across layers compared to vanilla self-improvement and other baselines, where middle layers undergo disproportionately larger updates than early and late layers. IMM model therefore brings better generalized reasoning capability.
Datasets GSM8K MATH SAT_Math MAWPS
[…]
Figure 13: ID performance with different k for scaling up test-time-computing Pass@k on MATH.
B.10 Complexity
Let $n$ be the number of model parameters, $T$ be the number of IMM iterations, $F(n)$ be the cost of one SFT training session. We calculate the complexity for IMM in the Table 10. The overall complexity is $\mathcal{O}(T\cdot F(n))$. Since fine-tuning dominates, especially for large models, the primary bottleneck is still the repeated SFT stages. Therefore, IMM introduce linear complexity on $n$, which can be overlooked compared with $\mathcal{O}(F(n))$, ensuring the scalability.
Appendix C Additional Discussion and Clarification
C.1 A Bitter Lesson: Not All LLMs Can Self-improve
During our experiment, we also find that not all the LLMs can self-improve on reasoning tasks. If LLM’s performance decreases after SFT, then our method may not let the merged model have a better performance compared with the original model and the model after SFT. This usually happens when the original model already has a good performance (reasoning ability), and learned reasoning ability can’t offset the generalization loss.
C.2 Why Importance-Weight Change Mismatch Happens?
We conclude two possible contributing factors to this observation: (i) Characteristics of SFT on Pre-trained LMs: Prior studies Merchant et al. (2020); Mosbach et al. (2020); Zhou and Srikumar (2021) have shown that during SFT, the early and late layers of pre-trained language models tend to undergo minimal changes. In particular, the late layers often preserve their original representations, suggesting a structural bias of SFT toward updating the middle layers. (ii) Inhibitory Effect of Self-improvement on Reasoning-critical Layers: As shown in Figure 9, models fine-tuned on real data exhibit more weight change in reasoning-critical layers (early and late layers) compared to those fine-tuned on self-synthesized data. In contrast, the middle layers show comparable levels of weight change in both settings. This indicates that the self-improvement process inherently inhibits updates to reasoning-critical layers, leading to disproportionate changes in the middle layers.
We further explain why middle layers contribute less to complex reasoning tasks. Prior work Li et al. (2024) shows that weaker, implicit reasoning signals tend to surface in the middle layers, whereas stronger, explicit reasoning—such as chain-of-thought reasoning—emerges primarily in the late (and occasionally early) layers. In our study, to solve complex reasoning tasks model generated long CoT reasoning path, which depends on late layers
In summary, superficial self-improvement leads to overfitting on middle-layer representations where weaker, implicit reasoning resides, due to both the inherent bias of SFT and self-generated data. In contrast, reasoning-critical layers, responsible for explicit CoT reasoning, remain largely unchanged, limiting the model’s ability to improve on more complex reasoning tasks. Actually, it’s very common that different layer has different behaviors for tasks Qing et al. (2024), and our work emphasizes it in reasoning.
Operation Complexity
SFT $\mathcal{O}(F(n))$
Compute $\delta^{t}$ $\mathcal{O}(n)$
Masking, scaling $\mathcal{O}(n)$
Merge update $\mathcal{O}(n)$
Overall Complexity $\mathcal{O}(T\cdot(F(n)+n))\approx\mathcal{O}(T\cdot F(n))$
Table 10: Time complexity of IMM update steps.
C.3 The Connection to Catastrophic Forgetting
@@
C.4 The Definition for Layers
We do not provide a rigorous theoretical definition or external citation for the terms "reasoning-trivial layers" and "reasoning-trivial layers". In our paper, we adopt a relative and empirical definition: "reasoning-trivial layers" refer to the layers that exhibit lower importance scores in comparison to others, and "reasoning-trivial layers" refer to the layers that exhibit higher importance scores based on our layer-wise reasoning importance analysis. While not formally defined, this relative notion is sufficient for our purposes. It allows us to identify and analyze the mismatch between reasoning-critical layers (i.e., those with high importance scores) and the layers undergoing the most weight change during self-improvement. This mismatch is central to our discovery of the superficial self-improvement phenomenon.
C.5 Why This Importance Score
We would like to clarify that while the identification of key layers has been widely explored in prior work, such as in model analysis, pruning Liu et al. (2025), and importance-based selection, our study does not aim to introduce a theoretical advancement in key layer selection itself. Rather, our contribution lies in uncovering a novel phenomenon: a mismatch between reasoning-critical layers and the layers experiencing the most weight change during self-improvement. We believe this observation offers a new perspective on how generalized reasoning capabilities may be hindered by superficial self-improvement. Building on this insight, we propose IMM as a method to mitigate this issue and improve the model’s generalization in reasoning tasks.
Compared to other popular evaluation such as gradient change, the metrics defined in Eq. (3) and Eq. (4) are more suitable for the type of analysis conducted in this work. Specifically, Eq. (3) directly measures "how much the parameters of a given layer have actually changed from the beginning to the end of training." This provides a clearer indication of how much information is retained or adjusted through the self-improvement process, which is more aligned with our goal of understanding where learning occurs across the model. In contrast, gradient change is more appropriate for analyzing how quickly or at which stage the model learns during training. We appreciate the suggestion and agree that gradient analysis can provide complementary insights. We will include gradient tracking in manuscript to help monitor training stability and to identify potential issues such as exploding or vanishing gradients during self-improvement cycles.
C.6 SFT Overfitting
Overfitting is a common issue in supervised learning. Indeed, our work does not dispute or repeat this general principle. Rather, our contribution lies in highlighting a previously underexplored phenomenon: in particular, when applied to reasoning tasks, self-improvement via SFT tends to exacerbate generalization degradation more severely than standard SFT on curated or distilled data.
This distinction is central to our study. While overfitting is ubiquitous in supervised learning, our empirical results demonstrate that SFT using self-generated data amplifies this risk, leading to a sharper decline in reasoning generalization and more pronounced model collapse. This opinon is also supported by our interpretability analysis. To the best of our knowledge, this amplification effect of self-improvement has not been explicitly analyzed in prior work. To support this claim, we conducted systematic experiments beyond the self-improvement setting in Section 5.4 and Appendix B.1. We also demonstrate that even in general SFT settings like distillation, our proposed method still improves reasoning generalization, highlighting the broader applicability and potential of our approach.
C.7 Model Scale and Overfitting
While it is intuitive to assume that smaller models are more susceptible to overfitting, prior work Dohmatob et al. (2025) suggests that overfitting severity depends non-monotonically on model size. Specifically, they report that models below a certain size threshold may exhibit pronounced overfitting, whereas models above that threshold may also overfit due to reduced margin for improvement on certain tasks.
In our study (Section 8), we scale model size to acknowledge this complexity and do not assert a universal trend. Our empirical results in Figure 6 and Figure 7 show inconsistent overfitting behavior across different model sizes and datasets, making it difficult to isolate model size as the primary factor behind model collapse. In fact, our cross-model and cross-dataset comparison suggests an interesting trend: when a model already performs strongly on a dataset, further supervised fine-tuning on self-generated (and correct) samples may inadvertently trigger overfitting due to reduced room for meaningful generalization.
C.8 How Other Methods Lead to Superficial Reasoning Learning
While existing approaches like data mixture Shumailov et al. (2023), data accumulation Gerstgrasser et al. (2024), adding real data Dohmatob et al. (2025), and data selection Guo et al. (2024b), have indeed been effective in expanding the distribution or sampling diversity, their application has primarily targeted non-reasoning tasks, for example, language modeling, open-ended generation, and summarization. These tasks are less sensitive to distributional sparsity and often benefit from simple accumulation or semantic filtering strategies that enhance diversity.
However, our work demonstrates that reasoning tasks are more vulnerable to model collapse due to complex reasoning tasks require structured, compositional, and often multi-hop inference capabilities. In such tasks, simply expanding diversity in a linguistic or syntactic manner often filters out complete and structured reasoning trajectories or repeat low quality data, with models failing to learn high-quality reasoning patterns with generalization. This is a critical gap that previous literature has not examined, especially in SFT reasoning data. This work diagnoses and mitigates model collapse for reasoning tasks under SFT, showing that naive distribution expansion can be ineffective or even worse. Our model interpretability analysis shows that using self-generated data will strengthen the harmful model weights updating. Therefore, data diversity method still hard to avoid it. In contrast, we propose that model merging provides a more robust and effective solution inspired by interpretability analysis. Empirically, we show that this strategy outperforms data-centric diversity enhancements in preserving general reasoning capability. This distinction between task types (reasoning vs. non-reasoning), model interpretability, and the limitations of prior methods is an important motivation of our work.
C.9 Related Works on LLM Reasoning
LLMs have demonstrated remarkable success across various reasoning tasks, including mathematical problem-solving, code generation, multi-modality, agent, and common-sense reasoning Yu et al. (2024b); Wang et al. (2023); Shi et al. (2023); Diao et al. (2025); Fu et al. (2024b); Wu et al. (2025); Shi et al. (2025); Jin et al. (2025); Ye et al. (2025); Diao et al. (2025); Zhang et al. (2025a, b). Beyond leveraging sophisticated prompting techniques to enhance reasoning capabilities Kojima et al. (2022); Wei et al. (2022); Zheng et al. (2024); Yao et al. (2024), many methods focus on fine-tuning LLMs with reasoning datasets to create more robust reasoners Lu et al. (2024); Yu et al. (2024b). For instance, approaches like SI Huang et al. (2022), STaR Zelikman et al. (2022), V-STaR Hosseini et al. (2024), and rSTaR Qi et al. (2025) fine-tune LLMs on task-specific datasets or synthesize reasoning data tailored for corresponding tasks. In addition to training models to generate correct answers, some studies introduce external verifiers Cobbe et al. (2021); Lightman et al. (2023); Hosseini et al. (2024); Yuan et al. (2024) that select the best answer from a set of candidate solutions.
Appendix D Potential Risks
Enhancing LLMs with self-improving generalized reasoning capability may introduce risks of unintended capability emergence, including misuse in adversarial contexts such as misinformation or manipulation. As the model gains broader reasoning abilities across domains, it may be used for enabling harmful applications with enhanced reasoning capability. This highlights the importance of pairing IMM with safe evaluation and alignment to ensure safe and responsible deployment.